40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 15 00:40:19 2019, gautam, Update, SUS, ETMY diagnosis plan ETMY_sensorSpectra.pdf
    Reply  Wed May 15 10:57:57 2019, gautam, Update, SUS, EY chamber opened 
       Reply  Wed May 15 17:46:24 2019, gautam, Update, SUS, ETMY inspection firstContactFiber.JPGETMY_sensorSpectra.pdf
          Reply  Wed May 15 19:36:29 2019, Koji, Update, SUS, ETMY instepction 
             Reply  Thu May 16 13:07:14 2019, gautam, Update, SUS, First contact residue removal etmy_sensors.pdfetmy_BRmode.pdf
                Reply  Thu May 16 23:31:55 2019, gautam, Update, SUS, ETMY suspension characterization 6x
                   Reply  Fri May 17 17:01:08 2019, gautam, Update, SUS, ETMY suspension characterization etmy_oplevs.pdf
          Reply  Mon May 20 11:33:46 2019, gautam, Update, SUS, ITMY inspection 
             Reply  Mon May 20 17:12:57 2019, gautam, Update, SUS, ETMY LL adjustment 
                Reply  Mon May 20 22:06:07 2019, gautam, Update, SUS, ITMY also kicked 
                   Reply  Tue May 21 00:15:21 2019, gautam, Update, SUS, Main objectives of vent achieved (?) etmy_oplevs_20190520.pdfETMY_cplxTF.pdfETMY_diagComp.pdfScreen_Shot_2019-05-21_at_12.37.08_AM.png
                      Reply  Tue May 21 21:33:27 2019, gautam, Update, SUS, ETMY HR face cleaned 
                         Reply  Wed May 22 11:53:50 2019, gautam, Update, SUS, ETMY EQ stops backed out ETMY_eigenmodes.pdf
Message ID: 14628     Entry time: Tue May 21 00:15:21 2019     In reply to: 14627     Reply to this: 14629
Author: gautam 
Type: Update 
Category: SUS 
Subject: Main objectives of vent achieved (?) 

Summary:

  1. ETMY now shows four suspension eigenmodes, with sensible phasing between signals for the angular DoFs. However, the eigenfrequencies have shifted by ~10% compared to 16 May 2019.
  2. PIT and YAW for ETMY as witnessed by the Oplev are now much better separated.
  3. ITMY can have its bias voltage set to zero and back to nominal alignment without it getting stuck.
  4. The sensing matrix for ETMY that I get doesn't make much sense to me. Nevertheless, the optic damps even with the "naive" input matrix.

So the primary vent objectives have been achieved, I think. 


Details:

  1. ETMY free-swinging data after adjusting LL and SIDE coils such that these were closer to half-light values
    • Attachment #1 - oplev witnessing the angular motion of the optic. PIT and YAW are well decoupled.
    • Attachment #2 - complex TF between the suspension coils. There is still considerable imbalance between coils, but at least the phasing of the signals make sense for PIT and YAW now.
    • Attachment #3 - DoFs sensed using the naive and optimized sensing matrices.
    • Attachment #4 - sensing matrix that the free swinging data tells me to implement. If the local damping works with the naive input matrix but we get better diagonality in the actuation matrix, I think we may as well stick to the naive input matrix.
  2. BR mode coupling minimization:
    • As alluded to in my previous elog, I tried to reduce the bounce mode coupling into the shadow sensor by rotating the OSEM in its holder.
    • However, I saw negligible change in the coupling, even going through a full pi radian rotation. I imagine the coupling will change smoothly so we should have seen some change in one of the ~15 positions I sampled in between, but I saw none.
    • The anomalously high coupling of the bounce mode to the shadow sensor readout is telling us something - I'm just not sure what yet.
  3. ITMY:
    • The offender was the LL OSEM, whose rotational orientation was causing the magnet to get stuck to the teflon part of the OSEM coil when the bias voltage was changed by a sufficiently large amount.
    • I rectified this (required adjustment of all 5 OSEMs to get everything back to half light again).
    • After this, I was able to zero the bias voltage to the PIT/YAW DoFs and not have the optic get stuck - huzzah 😀 
    • While I have the chance, I'm collecting the free-swinging data to see what kind of sensing matrix this optic yields.

Tomorrow and later this week:

  1. Prepare ETMY for first contact cleaning to remove the residual piece. 
    • Drag wipe the HR surface with dehydrated acetone 
    • Apply F.C. as usual, inspect the HR face after peeling for improvement if any.
    • This will give us a chance to practise the F.C.ing with the optic EQ-stopped (moving cage etc).
  2. Confirm ETMY actuation makes sense.
    • Use the green beam for an ASS proxy implementation?
  3. High quality close out pictures of OSEMs and general chamber layout.
  4. Anything else? Any other tests we can do to convince ourselves the suspensions are well-behaved?

While we have the chance:

  1. Fix the IPANG alignment? Because the TT drift/hysteresis problem is still of unknown cause.
  2. Check that the AS beam is centered on OMs 1-6?
  3. Recover the 70% AS light that is being diverted to the OMC?

Unrelated to this work: megatron is responding to ping but isn't ssh-able. I also noticed earlier to day that the IMC autolocker blinky wasn't blinking. So it probably requries a hard reboot. I left the lab for tonight so I'll reboot it tomorrow, but no nds data access in the meantime... 

Attachment 1: etmy_oplevs_20190520.pdf  34 kB  Uploaded Tue May 21 01:35:15 2019  | Hide | Hide all
etmy_oplevs_20190520.pdf
Attachment 2: ETMY_cplxTF.pdf  367 kB  Uploaded Tue May 21 01:36:28 2019  | Hide | Hide all
ETMY_cplxTF.pdf
Attachment 3: ETMY_diagComp.pdf  67 kB  Uploaded Tue May 21 01:36:50 2019  | Hide | Hide all
ETMY_diagComp.pdf
Attachment 4: Screen_Shot_2019-05-21_at_12.37.08_AM.png  60 kB  Uploaded Tue May 21 01:37:22 2019  | Hide | Hide all
Screen_Shot_2019-05-21_at_12.37.08_AM.png
ELOG V3.1.3-